OpenStudy (anonymous):

Simplify the following expression. a + 5c − b − c + a − b

OpenStudy (anonymous):

you can add or subtract like variables.

OpenStudy (anonymous):

what ^^ .....

OpenStudy (anonymous):

combine like terms, a+a= ? 5c-c= ? -b-b= ?

OpenStudy (anonymous):

i get that but i always get the signs mixed up and end up getting the problem wrong.

OpenStudy (anonymous):

i do that too sometimes lol i'll walk you threw it a+a= ? they are both positive so what you do is add the coefficients together, which in this case is 1 and 1+1 is 2, so you have 2a

OpenStudy (anonymous):

okay , so then would 5c-c= 4c?

OpenStudy (anonymous):

yeah then -b-b

OpenStudy (anonymous):

which would = -2b?

OpenStudy (anonymous):

exactly (: then put it all together

OpenStudy (anonymous):

2a+ -2b + 4c?

OpenStudy (anonymous):

wait, 2a - 2b + 4c?

OpenStudy (anonymous):

yepp thats it (:

OpenStudy (anonymous):

thank you! <#

OpenStudy (anonymous):

no problem (: